# Quantifying the Cross-sectoral Intersecting Discrepancies within Multiple Groups Using Latent Class Analysis Towards Fairness

### v2

@@ Abstract
The growing interest in fair AI development is evident. The “Leave No One Behind” initiative urges us to address multiple and intersecting forms of inequality in accessing services, resources, and opportunities, emphasising the significance of fairness in AI. This is particularly relevant as an increasing number of AI tools are applied to decision-making processes, such as resource allocation and service scheme development, across various sectors such as health, energy, and housing. Therefore, exploring joint inequalities in these sectors is significant and valuable for thoroughly understanding overall inequality and unfairness. This research introduces an innovative approach to quantify cross-sectoral intersecting discrepancies among user-defined groups using latent class analysis. These discrepancies can be used to approximate inequality and provide valuable insights to fairness issues. We validate our approach using both proprietary and public datasets, including EVENS and Census 2021 (England & Wales) datasets, to examine cross-sectoral intersecting discrepancies among different ethnic groups. We also verify the reliability of the quantified discrepancy by conducting a correlation analysis with a government public metric. Our findings reveal significant discrepancies between minority ethnic groups, highlighting the need for targeted interventions in real-world AI applications. Additionally, we demonstrate how the proposed approach can be used to provide insights into the fairness of machine learning.
1 Introduction
As AI systems become increasingly prevalent, ensuring fairness in their design and implementation is crucial [18]. AI fairness research spans various sectors, including healthcare [4, 3, 2], finance [33], and education [8]. However, studies on cross-sectoral intersecting fairness in AI are limited. Stiglitz [27] highlights that economic inequalities extend beyond income and wealth, affecting sectors such as education and health, implying correlated inequalities. At the same time, the “Leave No One Behind” (LNOB) principle urges addressing multiple, intersecting inequalities that harm individuals’ rights [29]. Therefore, we believe that research exploring joint inequalities across sectors is necessary to comprehensively understand overall inequality and gain insights into unfairness.
@@ 1 Introduction
Bias, Inequality, Fairness Many obstacles to accessing services, resources, and opportunities result from discriminatory laws, policies, and societal norms that marginalise specific groups [29]. Recent studies highlight concerns that AI-supported decision-making systems may also be influenced by biases [16], which can unfairly impact vulnerable groups, such as ethnic minorities, emphasising the need for research on AI system fairness [20]. A primary obstacle in advancing and implementing fair AI systems is the presence of bias [9]. In AI, bias can originate from various sources, including data collection, algorithmic design, and user interaction, as illustrated in Figure 2.
@@
Most AI systems heavily rely on data for their operations. This close connection means that any biases in the training data can be embedded into the algorithms, leading to biased predictions. Even if the data itself is not inherently biased, algorithms can still exhibit biased behaviour due to inappropriate design choices. These biased outcomes can influence AI systems in real-world applications, creating a feedback loop where biased data from user interactions further trains and reinforces biased algorithms, resulting in a vicious cycle [18].
From our perspective, bias stemming from data plays a crucial role in achieving fairness, as it can trigger a cascade of other biases, exacerbating fairness issues. Additionally, data bias may reflect various forms of social inequality in the real world. Inequality refers to disparities in opportunities and rewards based on different social positions or statuses within a group or society [24]. Inequality is considered unfair due to its causes, such as discrimination or failures in providing equal opportunities, and its consequences, such as resulting in objectionable disparities in status or power [24]. As previously mentioned, cross-sectoral intersecting inequality is crucial. Therefore, quantifying these disparities within the data is essential for understanding inequality and bias, ultimately leading to the development of fairer AI systems.
Discrepancy According to the LNOB principle of equal opportunities, everyone should ideally have equal access to public services and resources without discrepancies. We prefer the term “discrepancy” over “disparity” or “difference” because it suggests an unexpected difference. Discrepancies in data may indicate unfairness, biases, and inequalities, given that individuals should ideally be treated equally. It is worth noting that bias can arise from various inequalities, though not all inequalities are biases [24]. However, in the context of LNOB, for this study, “discrepancy” is more appropriate when discussing equal opportunity, as it can be seen as a specific type of inequality causing unfairness.
@@
Figure 2 illustrates the correlations between fairness, bias, inequality, and discrepancy. Essentially, fairness is indirectly linked with discrepancy, and discrepancy can contribute to unfairness. The difference between fairness and bias is that the former can be viewed as a technical issue, while the latter can be viewed as a social and ethical issue [9]. Furthermore, bias is a problem caused by historical and current social inequality [6], and inequality can manifest as discrepancies. Figure 2 starts with discrepancy and moves through inequality and bias to fairness. Therefore, our research focuses on quantifying cross-sectoral intersecting discrepancies between different groups, uncovering insights or patterns related to inequality, bias, and AI fairness.
@@
Background and Motivation Currently, there is limited research focusing on quantifying discrepancies. Most recent research on quantifying bias and/or inequality primarily revolves around resource allocation strategies and generally relies on objective data (e.g., [32]). However, these approaches have limitations and face challenges in effectively assessing and measuring biases in datasets unrelated to resource allocation. For example, in social science, much data is collected through questionnaires, which often include binary, categorical, or ordinal data types related to subjective responses and user experiences. These questionnaires may cover various aspects, resulting in intersecting, cross-sectoral, and high-dimensional data. Analysing data based on no more than two dimensions or sectors may overlook important information or patterns. Therefore, we believe that quantifying cross-sectoral intersecting discrepancies is valuable, as it can provide comprehensive insights and joint information.
@@
In our preliminary research (see Appendix A) in England, based on a Anonymous project survey question regarding health and digital services  [^2] Question 21: Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?, we observed a notable discrepancy in the Chinese group: 30.16% lacked English proficiency and 26.98% struggled to use the online system, while most other ethnic groups reported no concerns. These discrepancies, stemming from cultural differences, user experiences, or discrimination, contribute to inequality and may affect AI fairness. For instance, an AI system might wrongly assume Chinese individuals need more English support, leaving other ME groups with less assistance. Ideally, the percentages should be similar across ethnic groups in the LNOB context, indicating equal treatment or experience. However, investigating multiple and cross-sectoral questions simultaneously is challenging. For instance, the Anonymous project’s data contains similar questions for the energy and housing sectors, and current methods struggle to analyse these sectors jointly. Therefore, we propose an approach to quantify intersecting and cross-sectoral discrepancies for multiple ethnic groups by leveraging latent class analysis (LCA) [19].
LCA is a popular method in social science [5] because it identifies latent groups within a population based on observed characteristics or behaviours. LCA offers a flexible framework for exploring social phenomena and integrating with other analytical techniques. In this research, we use LCA to cluster intersecting and cross-sectoral data, encompassing questions across the health, energy, and housing sectors. This method enables us to derive latent classes and outcomes, with each class describing a distinct cross-sectoral user profile. This approach moves beyond defining user classes solely based on individual questions. More details of our proposed approach are presented in Section 2, with experiments and results reported in Section 4.
@@
2 Quantifying the Cross-sectoral Intersecting Discrepancies
The overall workflow of the proposed approach is shown in Figure 3, using a binary-encoded survey data as an example. It is noted that our approach is not limited to this specific format and can be applied to a wide range of similar problems. We will present more experiments with other datasets in Section 4 to further validate and showcase the features of our approach.
In Figure 3, Stage (1) shows the binary-encoded data $D$, where $\{Q_{1},Q_{2},...\}$ represents the selected questions from the survey data, covering user experiences across sectors. Meanwhile, $\{u_{1},u_{2},u_{3},...\}$ represents the collection of survey respondents. Here, $q$ represents the response options, for example, $q_{2,1}$ denotes the first option for $Q_{2}$. A ‘1‘ indicates a selected option, while ‘0’ indicates it was not selected. For other datasets, the encoding method should be chosen based on the data format and type. The blue arrow in Figure 3 illustrates the LCA process  [^3] StepMix (https://stepmix.readthedocs.io/en/latest/index.html) Python repository is used to implement LCA in this research., which includes hyperparameter selection and model fitting.
The advantage of using LCA is straightforward: it considers the joint probability distribution of all variables. This means potential inequalities or discrepancies can be analysed jointly. Once we obtain the distributions of latent classes $\{c_{1},c_{2},c_{3},...\}$ over user-defined groups $\{e_{1},e_{2},e_{3},...\}$ (as shown in Figure 3 Stage (2)), we can calculate the discrepancy $\Delta$.
@@ 2 Quantifying the Cross-sectoral Intersecting Discrepancies
14:Output: Discrepancy matrix $S$ with size $|G|\times|G|$
@@
To initiate the quantification process, the proportions $r$ of samples from each user-defined group within each latent class need to be calculated, as detailed in Stage (3) in Figure 3. This calculation can be performed using $r_{c}^{e}=N^{e}_{c}/N^{e},\forall e\in G,c\in C$. The reason for calculating $r$ is that user-defined groups may have different numbers of samples; therefore, using percentages for subsequent analyses ensures fairness and consistency.
Subsequently, we can derive a matrix of results characterised by dimensions $|G|\times|C|$, as shown in Figure 3 (3). Within this matrix, each row corresponds to the proportions of samples from a specific group within each latent class. It is important to note that in this context, each latent class effectively represents an individual user profile and can be viewed as a distinctive feature. Consequently, each row within the matrix may be employed as a feature vector denoted as $R_{e}$, serving as a representation of a specific group within the feature space.
In the assessment of discrepancy between two vectors, various methods may be employed, including the Euclidean distance, Kullback-Leibler Divergence, Earth Mover’s Distance, and Manhattan Distance, among others. In our approach, we propose the utilisation of Cosine Similarity to calculate the discrepancy, which is defined as $\Delta=1-\cos(\theta)=1-\frac{\mathbf{A}\cdot\mathbf{B}}{\|\mathbf{A}\|\|\mathbf{B}\|}$. Finally, we can iteratively calculate $\Delta$ between any pairs of vectors $R$ and obtain the discrepancy matrix $S$ (as shown in Stage (4) of Figure 3). The AVG column in Stage (4) contains the mean discrepancy values for $e$, which can be viewed an approximation for how each $e$ is different from others.
We suggest the use of Cosine Similarity due to its inherent characteristics, including a natural value range spanning from 0 to 1. Importantly, it does not necessitate additional normalisation procedures. This metric, possessing with a fixed value range, enhances comparability and offers support to subsequent AI fairness research. The proposed approach is summarised in Algorithm 1.
3 Related Work
@@
4 Experiments
4.1 The Anonymous project
Within the Anonymous project444 More dataset details can be found in the Appendix A., a systematic online survey was conducted to explore the experiences of individuals from ME groups in the UK, focusing on digitalised health, energy, and housing aspects. To examine cross-sectoral intersecting discrepancies among the seven ethnic groups (as shown in Table 1), we selected three questions from the Anonymous survey data related to these sectors. The answers from ME participants formed the dataset for this experiment. For England and Scotland, we have 594 and 284 samples, respectively. Due to varying sample sizes across ethnicities, we calculated discrepancies separately for each region.
@@ 4.1 The Anonymous project
The discrepancy results for England are presented in the left part of Table 1. The table shows that the Chinese group has the largest average (AVG) discrepancy value compared to other groups. Meanwhile, the Indian group exhibits the smallest discrepancy with the Bangladeshi group and also shows similarity to the Pakistani group. This is likely due to their close geographical locations and similar cultural backgrounds and lifestyles.
The right part of Table 1 presents the results for Scotland, showing similar outcomes: the Chinese group is distinct from others, while the Bangladeshi group is similar to the Pakistani group. The differences between England and Scotland may be attributed to their different policies and circumstances. We hypothesise that the primary reason for the Chinese group standing out is a lack of English proficiency. This is supported by our preliminary research mentioned in Section 1, which shows a significant number of Chinese participants expressing this concern. Additionally, BBC News [23] reports that the Chinese community experiences some of the highest rates of racism among all ethnic groups in the UK. Our discrepancy values may help explain this, as the Chinese group shows different experiences in digitalised online services.
Table 1: The matrix of discrepancies between 7 ethnic groups for Anonymous project’s England and Scotland data. The AVG denotes the average discrepancy value for one group.
@@
4.2 EVENS

### v3

@@ Abstract
The growing interest in fair AI development is evident. The “Leave No One Behind” initiative urges us to address multiple and intersecting forms of inequality in accessing services, resources, and opportunities, emphasising the significance of fairness in AI. This is particularly relevant as an increasing number of AI tools are applied to decision-making processes, such as resource allocation and service scheme development, across various sectors such as health, energy, and housing. Therefore, exploring joint inequalities in these sectors is significant and valuable for thoroughly understanding overall inequality and unfairness. This research introduces an innovative approach to quantify cross-sectoral intersecting discrepancies among user-defined groups using latent class analysis. These discrepancies can be used to approximate inequality and provide valuable insights to fairness issues. We validate our approach using both proprietary and public datasets, including both EVENS and Census 2021 (England & Wales) datasets, to examine cross-sectoral intersecting discrepancies among different ethnic groups. We also verify the reliability of the quantified discrepancy by conducting a correlation analysis with a government public metric. Our findings reveal significant discrepancies both among minority ethnic groups and between minority ethnic groups and non-minority ethnic groups, emphasising the need for targeted interventions in policy-making processes. Furthermore, we demonstrate how the proposed approach can provide valuable insights into ensuring fairness in machine learning systems.
Index Terms:
discrepancy, machine learning, fairness
I Introduction
[…]
The “Leave No One Behind” (LNOB) principle emphasises the importance of addressing multiple, intersecting inequalities that harm individuals’ rights [1]. Intersecting inequality refers to the compounded disadvantages that arise from both the overlap of marginalised social categories (e.g., being female and living with a disability) and the intersection of multiple, mutually reinforcing dimensions of exclusion (e.g., deprivation in both health and education) [2]. Meanwhile, the increasing adoption of AI tools in decision-making processes across various sectors, including health [3], energy [4], and housing [5], underscores the urgency of ensuring fairness in their design and implementation [6], as unfair AI systems may risk deepening existing inequalities.
AI fairness research spans various sectors, including healthcare [7, 8, 9], finance [10], and education [11]. However, research on cross-sectoral intersecting AI fairness remains limited. The term “cross-sectoral intersecting” refers to the interaction and overlap of multiple sectors, such as healthcare, housing, and energy.
To address this gap, we propose quantifying cross-sectoral intersecting discrepancies between groups. These discrepancies refer to differences in user profiles across groups and can serve as a proxy for underlying inequalities, providing valuable insights for stakeholders. Additionally, the quantified discrepancies in data can offer insights into AI fairness when the data is used to train models. According to the LNOB principle of equal opportunities, everyone should ideally have equal access to public services and resources without discrepancies. We prefer the term “discrepancy” over “disparity” or “difference” because it suggests an unexpected difference.
@@ I Introduction
Bias and Fairness Recent studies highlight concerns that AI-supported decision-making systems may be influenced by biases [12], which can unfairly impact vulnerable groups, such as ethnic minorities, emphasising the need for research on AI system fairness [13]. A primary obstacle in advancing and implementing fair AI systems is the presence of bias [14]. In AI, bias can originate from various sources, including data collection, algorithmic design, and user interaction, as illustrated in Fig. 1.
@@
In particular, most AI systems rely on data for training and prediction. This close connection means that any inherent biases in the training data can be propagated and embedded into the AI systems, leading to biased predictions, the so called bias-in and bias-out. Even if the data itself is not inherently biased, algorithms can still exhibit biased behaviour due to inappropriate design and configuration choices. These biased outcomes can influence AI systems in real-world applications, creating a feedback loop where biased data from user interactions further trains and reinforces biased algorithms, resulting in a vicious cycle [6].
Bias stemming from data is a crucial factor affecting fairness, as inappropriate handling of it may trigger a cascade of other biases, exacerbating fairness issues.
@@
Discrepancy in data may indicate inequalities and lead to biases and unfairness, given that individuals should ideally be treated equally in ideal situations. Fig. 2 illustrates the correlations between fairness, bias, inequality, and discrepancy. Essentially, fairness is indirectly linked with discrepancy, and discrepancy can contribute to unfairness. The difference between fairness and bias is that the former can be viewed as a technical issue, while the latter can be viewed as a social and ethical issue [14]. Furthermore, bias is a problem caused by historical and current social inequality [15], and inequality can manifest as discrepancies. Fig. 2 starts with discrepancy and moves through inequality and bias to fairness. Therefore, our research focuses on quantifying cross-sectoral intersecting discrepancies among different groups, with an aim to uncover insights or patterns related to inequality, bias, and AI fairness.
@@
Background and Motivation Currently, there is limited research focusing on quantifying discrepancies. Most recent research on quantifying bias and/or inequality primarily revolves around resource allocation strategies and generally relies on objective data (e.g., [16]). However, these approaches have limitations and face challenges in effectively assessing and measuring bias or discrepancy in datasets unrelated to resource allocation. For example, in social sciences, much data is collected through questionnaires, which often include binary, categorical, or ordinal data types related to subjective responses and user experiences. These questionnaires may cover various aspects, resulting in intersecting and cross-sectoral data of high dimensions. Analysing data based on no more than two dimensions or sectors may overlook important information or patterns. Therefore, we believe that quantifying cross-sectoral intersecting discrepancies is valuable, as it can provide comprehensive insights.
@@
In our preliminary research (see Appendix A) in England, based on a Anonymous project survey question regarding health and digital services [^2] Question 21: Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?, we observed a notable discrepancy in the Chinese group: 30.16% lacked English proficiency and 26.98% struggled to use the online system, while most other ethnic groups reported less or no concerns. These discrepancies, stemming from cultural differences, user experiences, or discrimination, contribute to inequality and may affect AI fairness. For instance, an AI system might inappropriately assume that only Chinese individuals require English language support, thereby neglecting other ME groups who may also need assistance.
However, investigating multiple and cross-sectoral questions simultaneously is challenging. For instance, the Anonymous project’s data contains similar questions for the energy and housing sectors, and current methods struggle to analyse these sectors jointly. Therefore, we propose an approach to quantify intersecting and cross-sectoral discrepancies for multiple ethnic groups by leveraging latent class analysis (LCA) [17].
LCA is a popular method in social science [18] because it can identify latent groups within a population based on observed characteristics or behaviours. LCA offers a flexible framework for exploring social phenomena and integrating with other analytical techniques. In this research, we use LCA to cluster intersecting and cross-sectoral data, encompassing questions across the health, energy, and housing sectors. This method enables us to derive latent classes and outcomes, with each class describing a distinct cross-sectoral user profile. This approach moves beyond defining user classes solely based on individual questions. More details of our proposed approach are presented in Section II, with experiments and results reported in Section IV.
@@
II Quantifying the Cross-sectoral Intersecting Discrepancies
The overall workflow of the proposed approach is shown in Fig. 3, using a binary-encoded survey data as an example. It is noted that our approach is not limited to this specific format and can be applied to a wide range of similar problems. We will present more experiments with other datasets in Section IV to further validate and showcase the features of our approach.
In Fig. 3, Stage (1) illustrates the binary-encoded data $D$, where $\{Q_{1},Q_{2},\dots\}$ represents the selected survey questions, covering user experiences across different sectors. Similarly, $\{u_{1},u_{2},u_{3},\dots\}$ denotes the set of survey respondents. Here, $q$ represents the response options; for example, $q_{2,1}$ refers to the first option for $Q_{2}$. A value of ‘1’ indicates a selected option, while ‘0’ signifies that it was not selected. Each respondent’s responses can be represented as a vector $\mathbf{q}$. The set of all indicator variables is denoted by $X$, with $\mathbf{q}\in\mathcal{X}$, where $\mathcal{X}$ is the space of all possible response vectors. For other datasets, the encoding method should be selected based on the data format and type. The red arrow in Fig. 3 illustrates the LCA process [^3] StepMix (https://stepmix.readthedocs.io/en/latest/index.html) Python repository is used to implement LCA in this research., which includes hyperparameter selection and model fitting.
For simplicity, the LCA process is defined in Equation 1, where $\theta_{C}$ and $\theta_{X}$ specify the marginal distribution of the latent classes $C$ and the class-conditional distribution of the indicator variables $X$, respectively. Here, $\theta=\left(\theta_{C},\theta_{X}\right)$, and each latent class is denoted by $c$. In fact, the LCA is specified by a set of $\theta\in\Theta$, where $\left(\theta_{C},\theta_{X}\right)\in\left(\Theta_{C}\times\Theta_{X}\right)$.
$p(c,\mathbf{q};\theta)=p\left(c;\theta_{C}\right)p\left(\mathbf{q}\mid c;\theta_{X}\right)$ (1)
The advantage of using LCA is straightforward: it considers the joint probability distribution of all variables. This means potential inequalities or discrepancies can be analysed jointly. Once we obtain the distributions of latent classes $\{c_{1},c_{2},c_{3},...\}$ over user-defined groups $\{e_{1},e_{2},e_{3},...\}$ (as shown in Fig. 3 Stage (2)), we can calculate the discrepancy $\Delta$.
@@ II Quantifying the Cross-sectoral Intersecting Discrepancies
14:Output: Discrepancy matrix $S$ of size $|G|\times|G|$
@@
To initiate the quantification process, the proportions $r$ of samples from each user-defined group within each latent class need to be calculated, as detailed in Line 6 in Algorithm 1. This calculation can be performed using $r_{c}^{e}=N^{e}_{c}/N^{e},\forall e\in G,c\in C$. The reason for calculating $r$ is that user-defined groups may have different numbers of samples; therefore, using percentages for subsequent analyses ensures fairness and consistency.
Subsequently, we can derive a matrix of results characterised by dimensions $|G|\times|C|$, as shown in Fig. 3 (3). Within this matrix, each row corresponds to the proportions of samples from a specific group within each latent class. It is important to note that in this context, each latent class effectively represents an individual user profile and can be viewed as a distinctive feature. Consequently, each row within the matrix may be employed as a feature vector denoted as $\mathbf{r}_{e}$, serving as a representation of a specific group within the feature space.
In the assessment of discrepancy between two feature vectors, various methods may be employed, including the Euclidean distance, Kullback-Leibler Divergence, Earth Mover’s Distance, and Manhattan Distance, among others. In our approach, we propose the utilisation of Cosine Similarity to calculate the discrepancy, which is defined as $\Delta=1-\cos(\theta)=1-\frac{\mathbf{r}_{e}\cdot\mathbf{r}_{{e}^{\prime}}}{\|\mathbf{r}_{e}\|\|\mathbf{r}_{e^{\prime}}\|}$. Finally, we can iteratively calculate $\Delta$ between any pairs of vectors $\mathbf{r}$ and obtain the discrepancy matrix $S$ (as shown in Stage (4) of Fig. 3). The AVG column in Stage (4) contains the mean discrepancy values for $e$, which can be viewed an approximation for how each $e$ is different from others.
We suggest the use of Cosine Similarity due to its inherent characteristics, including a natural value range spanning from 0 to 1 as $\mathbf{r}_{e}$ and $\mathbf{r}_{e^{\prime}}$ contain no negative values. Importantly, it does not necessitate additional normalisation procedures. This metric, possessing with a fixed value range, enhances comparability and offers support to subsequent AI fairness research. The proposed approach is summarised in Algorithm 1.
III Related Work
@@
IV Experiments
IV-A The Anonymous project
Within the Anonymous project444 More dataset details can be found in the Appendix A., a systematic online survey was conducted to explore the experiences of individuals from ME groups in the UK, focusing on digitalised health, energy, and housing aspects. To examine cross-sectoral intersecting discrepancies among the seven ethnic groups (as shown in Table I), we selected three questions from the Anonymous survey data related to these sectors. The answers from ME participants formed the dataset for this experiment. For England and Scotland, we have 594 and 284 samples, respectively. Due to varying sample sizes across ethnicities, we calculated discrepancies separately for each region.
@@ IV-A The Anonymous project
The discrepancy results for England are presented in the left part of Table I. The table shows that the Chinese group has the largest average (AVG) discrepancy value compared to other groups. Meanwhile, the Indian group exhibits the smallest discrepancy with the Bangladeshi group and also shows similarity to the Pakistani group. This is likely due to their close geographical locations and similar cultural backgrounds and lifestyles.
The right part of Table I presents the results for Scotland, showing similar outcomes: the Chinese group is distinct from others, while the Bangladeshi group is similar to the Pakistani group. The differences between England and Scotland may be attributed to their different policies and circumstances. We hypothesise that the primary reason for the Chinese group standing out is a lack of English proficiency. This is supported by our preliminary research mentioned in Section I, which shows a significant number of Chinese participants expressing this concern. Additionally, BBC News [31] reports that the Chinese community experiences some of the highest rates of racism among all ethnic groups in the UK. Our discrepancy values may help explain this, as the Chinese group shows different experiences in digitalised online services.
To further verify the reliability of the discrepancy computation, we used PCA dimensionality reduction to visualise the relationships among ME groups based on the percentage distributions (as shown in Fig. 3 (3)) of each ME group across different latent classes. The patterns presented in Fig. 4 align with the findings obtained from the analysis of Table I (England). Overall, the Chinese group continues to stand out, positioned further away from other groups, indicating it has the highest average discrepancy. The Mixed and Bangladeshi groups are relatively distant from others, as they are scattered on opposite sides. Indian, Bangladeshi, and Pakistani groups are closer to each other compared to other groups, implying a similarity supported by the discrepancy values in Table I. Meanwhile, the African and Caribbean groups are the nearest to each other, as represented by the double arrows between them.
TABLE I: The matrix of discrepancies between 7 ethnic groups for Anonymous project’s England and Scotland data. The AVG denotes the average discrepancy value for one group.
@@
Figure 4: The PCA visualisation of relationships among ME groups based on percentage distributions across latent classes on PRIME dataset (England). The dashed line with an arrow indicates the distance from one dot to its nearest neighbour.
IV-B EVENS